Find the critical numbers of the function. p-4. p2 +1 p = (smaller value) p = (larger value)
Answer #1

d dp h(P) = P-4 p²4 P-4 h(P) = ene (h(p)) = ( Pop ) lo (P-4)).(p?+1) – (P-4) truth (P241) (p?+1)? (1-0) (p2+1) -(8-4) (22+0)